OpenStudy (anonymous):

Find the length and width of a rectangle that has an area of 128 square feet and whose perimeter is a minimum.

OpenStudy (anonymous):

\[perimeter \ge 32\sqrt{2}\]

OpenStudy (anonymous):

the options are.. l=16feet; w=8 feet l=32 feet; w=4 feet l=w=8 square root of 2 ( sorry cant find the square root sign)

OpenStudy (anonymous):

or..

OpenStudy (anonymous):

l=64 square root of 2 feet;; w= square root 2 feet or l=64 feet w=2 feet

OpenStudy (anonymous):

32,4

OpenStudy (anonymous):

thank you so much! would you be able to help me with this problem as well?

OpenStudy (anonymous):

On a given day, the flow rate F (cars per hour) on a congested roadway is given by F=v/28+0.07v^2 where v is the speed of the traffic in miles per hour. What speed will maximize the flow rate on the road? Round your answer to the nearest mile per hour.